OpenStudy (kimberly101):

Use equations to solve the following questions At a certain store, a football costs $15 more than a basketball. If 3 basketballs and 2 footballs cost a total of $115, find the cost of each ball.

OpenStudy (zzr0ck3r):

f = b + 15 3b+2f = 115

OpenStudy (zzr0ck3r):

you got it from there?

OpenStudy (zzr0ck3r):

-3b +3f = 45 3b + 2f = 115